1. ECN Inc. has a 5-year bond outstanding that has 7% coupon rate. If the appropriate discount rate for such a bond is 10%, what is the appropriate price for the semi-annual coupon paying bond? Assume a $1000 face value (par value).

2. The yield on a one-year bond is 6% today and is expected to be 8.5% next year. Based on the expectations theory, what is the yield of a two year bond today?
Answer #1

1. Answer is $813.07

C= 1000*7%/2=35

R= 10%/2= 5%

N = 10*2=20

FV= 1000

Price = (C*((1-(1+r)^-n)/r)+(FV/(1+r)^n) =(35*((1-(1.05^-20))/0.05))+(1000/(1.05^20))= $813.07

2. Answer is 15.01%

Yield = ((1.06)*(1.085)) - 1=1.1501-1=15.01%